Review | Constants Periodic Table Water ionizes by the equation H2 О(1) — н* (аq) + ОН (ад) The extent of the reaction is small in pure water and dilute aqueous solutions. This reaction creates the following relationship between [H ] and [OH ]: Kw 3D н'|Он ] Keep in mind that, like all equilibrium constants, the value of Kw changes with temperature
Part C At a certain temperature, the pH of a neutral solution is 7.26. What is the value of Kw at that temperature? Express your answer numerically using two significant figures View Available Hint(s) VOAD ? Kw = Submit
Answer #1

pH = 7.26

we have a relation between pH and [H3O+]

[H3O+] = 10-pH = 10-7.26

[H3O+] = 5.49 x 10-8 M

from Kw expression

it is clear

[H3O+] = [OH-] = 5.49 x 10-8

Kw = [H3O+][OH-]

Kw = [5.49 x 10-8][5.49 x 10-8]

Kw = 3.0 x 10-15